# RL2Grid: Benchmarking Reinforcement Learning in Power Grid Operations

### v1

@@ 5 Related Work
There have been several attempts to develop benchmarks for sequential decision-making in power system operations, but they often focus on smaller-scale problems and/or simplified setups (Chen et al.,, 2022). Examples include python-microgrid for simulating microgrids (Henri et al.,, 2020), CityLearn for demand response and urban energy management (Vazquez-Canteli et al.,, 2020), and gym-ANM for small electricity distribution networks (Henry and Ernst,, 2021). RL environments for electric vehicle (EV) charging and electricity markets have also been introduced (Zhang et al.,, 2020). Recently, SustainGym spanned diverse tasks ranging from EV charging to carbon-aware data center job scheduling (Yeh et al.,, 2023). The ARPA-E GO Competition provides a realistic, large-scale benchmark for power grid operations (ARPA-E,, 2023), but is more-so geared towards offline optimization approaches than online sequential decision-making. On the methodological side, recent contributions in the field include works on cascading failure mitigation, demand response optimization, and real-time grid control using RL (Matavalam et al.,, 2022; Lehna et al.,, 2023; van der Sar et al.,, 2024). Nonetheless, these works are more geared towards methodological advancements rather than proposing a benchmark. For this reason, we refer the reader to recent reviews for details on RL applications in power grid operations (Li et al.,, 2023; Su et al.,, 2024).
@@ 6 Tackling the Challenges of Power Grids with RL
Applying RL in power grids presents numerous open problems, each offering significant opportunities for advancing both grid operations and RL methodologies (Marot et al., 2022b, ). While we address a subset of these challenges via our benchmark, there remains ample room for future work.
6.1 RL Methodologies of Importance for Power Grids
Different RL techniques have the potential to be beneficial in addressing open grid problems. However, there are also potential risks—e.g., with respect to safety, reliability, and robustness—that are important to address. In the following, we summarize interesting avenues for future research.
Safe RL. Safety is paramount in grid operations. Safe RL methods aim to ensure that learning and control policies adhere to strict safety constraints, preventing actions that could lead to blackouts or equipment damage (Donnot,, 2020). Ensuring safety while optimizing performance is a critical area of research (Garcıa and Fernández,, 2015; Marzari et al.,, 2025). In particular, incorporating novel CMDP-based algorithms and formal methods can be particularly beneficial for ensuring that solutions adhere to physical and operational limits (Liu et al., 2021b, ; Marchesini et al.,, 2023).
@@ 6.1 RL Methodologies of Importance for Power Grids
Hierarchical control and multi-agent RL. Power grids operate across many hierarchical levels, from individual substations to entire regions. Effective coordination within and across these levels is crucial for maintaining efficient and reliable operations. Hierarchical RL methods can be developed to manage multi-level control tasks, in a way that addresses the scale and complexity of grid operations (Pateria et al.,, 2021; Aydeniz et al.,, 2023). Another promising direction is the use of multi-agent representations. Given the vast and distributed nature of power grids, scalability can be enhanced by dividing the grid into distinct agents, each responsible for its own operations. Multi-agent RL (MARL) frameworks can enable these agents to learn and coordinate actions (Papoudakis et al.,, 2021; Marchesini et al.,, 2024), to improve overall grid performance while managing local contingencies more effectively.
Robust RL. The integration of renewable energy sources introduces significant variability and uncertainty into power grids, leading to non-stationary environments. RL algorithms need to adapt to these evolving dynamics to ensure stable and efficient grid operations despite fluctuating supply and demand profiles. Handling non-stationarity and generalization are thus critical research directions (Moos et al.,, 2022; Marchesini and Amato,, 2023).
@@ 6.2 Improving Realism of Power Grid Environments
While RL2Grid aims to promote initial advancements in RL methodologies of relevance to power grids, it is important to acknowledge that this is only a first step. Notably, building on these advancements to develop “last-mile” deployed solutions will require further improvements in the realism of power grid environments. We highlight several important directions in this regard.
@@ 7 Conclusions
Power grids are essential in combating climate change, requiring a transition to low-carbon energy and enhanced resilience against climate-induced extremes. The integration of variable renewable energy sources introduces complexities and uncertainties in grid operations, posing significant challenges for human operators and traditional power system solvers. Our work aims to foster progress towards these challenges by introducing RL2Grid, a benchmark designed to bridge the gap between current grid management practices and methodological research in RL. RL2Grid provides a standardized interface for power grid environments, featuring common rewards, state spaces, action spaces, and safety constraints across a pre-designed set of diverse and complex grid tasks in order to provide a common ground for monitoring and promoting progress. We perform a comprehensive evaluation of the performance of DQN, PPO, LagrPPO, SAC, and TD3 on RL2Grid tasks, including versions augmented with domain-informed heuristics aimed at improving performance and sample efficiency, and find that there is still significant room for improvement in the performance of these methods. By offering a standardized platform for RL research in the context of power grids, RL2Grid aims to accelerate algorithmic innovation towards improving power grid operations amidst the evolving challenges posed by climate change.
@@ Appendix B RL baselines
PPO (Schulman et al.,, 2017) and Lagrangian PPO (LagrPPO) (Stooke et al.,, 2020). A PPO agent uses its neural network to directly approximate a policy. The agent learns the policy parameters by simplifying the TRPO (Schulman et al.,, 2015) algorithm, using a computationally tractable clipped objective. This clipping mechanism prevents large changes to the policy that could destabilize the training. At a high level, such a surrogate objective balances policy improvement and limits the divergence between policy updates. To drive the policy training, PPO also learns an advantage function to determine how much better (or worse) taking an action is compared to the expected value. By employing different probability distributions as a policy, a PPO agent can deal with both continuous (redispatching) and discrete (topological) actions. The Lagrangian version applies the same intuitions while learning additional value functions for each constraint. It then changes to policy training by considering the Lagrangian in Equation 2. In more detail, Lagrangian algorithms take gradient ascent steps in $\pi$ and descent steps in $\bm{\lambda}$ to trade off safety and task performance. These methods focus on satisfying the constraints using penalties $\bm{\lambda}$ that grows unbounded when constraints are violated. When constraints are satisfied, $\bm{\lambda}$ scale down (to zero), allowing the algorithm to maximize the task objective.
@@ Appendix D State Space
For the interested RL practitioner, we refer to the original Grid2Op documentation for exhaustive descriptions of these features (Donnot,, 2020).
@@ Appendix F Hyperparameters
Shared N° parallel envs 10 10
@@
$\alpha$ 0.1, 0.3, 0.6 0.1
$\beta$ 0.1, 0.3, 0.6 0.3
$\eta$ 0.1, 0.3, 0.6 0.3 - 0.6
$\omega$ 0.1, 0.3, 0.6 0.3
$\lambda$ 0, 50 0 (TLO), 50 (LSI)
DQN Train frequency 20, 100, 1000 20
Target network update 500, 1000, 10000 1000
Buffer size 100000, 250000, 500000, 1000000 500000
@@
Entropy regularization 0.2 0.2
Noise clip 0.5 0.5
@@
Buffer size 100000, 250000, 500000, 1000000 500000
@@ Appendix G Omitted Figures in Section 5
Figure G.1 shows the training curves for the (discrete) topological action spaces. Due to the strict time limit imposed on the computation nodes (see Section 4) and the different computational requirements of the algorithms, not all the baselines perform the same number of steps in the time limit. [^10] The demands and limited performance of the topological baselines led us to exclude the results with the complete action space (i.e., difficulty set to 4). Additionally, despite the grid search of Table F.1, some baselines achieved lower performance than expected (e.g., SAC and DQN in the bus14 scenarios). We will keep working on the benchmark to find better parameters, run longer experiments, and keep the following Figures updated.

### v2

@@ 5 Related Work
Recent attempts to develop sequential decision-making in power system operations benchmarks often focus on small-scale problems and/or simplified setups (Chen et al.,, 2022). Examples include Henri et al., (2020) for microgrids, CityLearn for demand response and urban energy management (Vazquez-Canteli et al.,, 2020), and gym-ANM for small electricity distribution networks (Henry and Ernst,, 2021). RL environments for electric vehicle (EV) charging and electricity markets have also been introduced (Zhang et al.,, 2020). Recently, SustainGym spanned diverse tasks ranging from EV charging to carbon-aware data center job scheduling (Yeh et al.,, 2023). The ARPA-E GO Competition provides a large-scale benchmark for power grid operations, but is more-so geared towards offline optimization approaches (ARPA-E,, 2023). On the methodological side, recent contributions in the field include works on cascading failure mitigation, demand response optimization, and real-time grid control using RL (Matavalam et al.,, 2022; Lehna et al.,, 2023; van der Sar et al.,, 2024). Nonetheless, these works are more geared towards methodological advancements rather than proposing a benchmark. For this reason, we refer the reader to recent reviews for details on RL applications in power grid operations (Li et al.,, 2023; Su et al.,, 2024).
Relationship with Grid2Op.
Grid2Op is an open-source power simulation framework designed by RTE France to model power grid controllers (RTE France,, 2020). It simulates grid operations, requiring that grids work for long horizons in a way that is robust to contingency events, as well as adhering to physical and operational constraints. For the latter, Grid2Op models: (i) cooldown periods to prevent immediate reconnection of disconnected lines, and limits on the frequency of actions on the same line to avoid asset degradation; (ii) limited thermal capacity of transmission lines; (iii) ramp rates on generators that restrict how much power generation can change between time periods; and (iv) adherence to AC power flow constraints. Grid2Op has been primarily used as the base for the L2RPN competitions (Marot et al., 2020a, ; Marot et al.,, 2021; Marot et al., 2022a, ). While RL methods have been used for L2RPN, they fail to provide a common ground to foster advancements in the field—each method uses custom input features and action spaces of (very) limited size, often without providing sufficient evidence on how and why these spaces were considered. Hence, to date, there is no standardized solution that allows RL researchers to easily get started in this field and compare over an established benchmark. [^9] Appendix A further discusses the relationship between L2RPN and RL2Grid.
@@ 6 Tackling the Challenges of Power Grids with RL
Applying RL in power grids presents numerous open problems, each offering significant opportunities for advancing both grid operations and RL methodologies (Marot et al., 2022b, ). While we address a subset of these challenges via our work, there remains ample room for future work.
6.1 Relevant RL Methodologies
RL has the potential to be beneficial in addressing open grid problems. There are also potential risks (e.g., with respect to safety, reliability, and robustness) that are important to address. Here, we summarize interesting avenues for future research.
Safe RL. Safety is paramount in power grid operations. Safe RL methods aim to ensure that learning and control policies adhere to strict safety constraints, preventing actions that could lead to blackouts or equipment damage (RTE France,, 2020). Ensuring safety while optimizing performance is a critical area of research (Garcıa and Fernández,, 2015; Marzari et al.,, 2025). In particular, incorporating novel algorithms based on our CMDP representations can be particularly beneficial for ensuring that solutions adhere to physical and operational limits (Liu et al.,, 2021; Marchesini et al.,, 2023).
@@ 6.1 Relevant RL Methodologies
Hierarchical control and multi-agent RL. Power grids operate across multiple hierarchical levels, from individual substations to entire regions. Effective coordination within and across these levels is crucial for maintaining efficient and reliable operations. Hierarchical RL methods can be developed to manage multi-level control tasks, in a way that addresses the scale and complexity of grid operations (Pateria et al.,, 2021; Aydeniz et al.,, 2023). Another promising direction is the use of multi-agent representations. Given the vast and distributed nature of power grids, scalability can be enhanced by dividing the grid into distinct areas or agents, each responsible for its own operations. Multi-agent RL (MARL) frameworks can enable these agents to learn and coordinate actions (Papoudakis et al.,, 2021; Marchesini et al.,, 2024), to improve overall grid performance while managing local contingencies more effectively.
Robust RL. The integration of renewable energy sources introduces significant variability and uncertainty into power grids, leading to non-stationary environments. RL algorithms need to adapt to these evolving dynamics to ensure stable and efficient grid operations despite fluctuating supply and demand profiles. Handling non-stationarity is thus a critical research direction (Moos et al.,, 2022; Marchesini and Amato,, 2023).
@@ 6.2 Improving Realism of Power Grid Environments
It is important to acknowledge that RL2Grid is only a first step. Notably, developing “last-mile” deployable solutions will require further improvements in the realism of power grid environments, which we now discuss.
@@ 7 Conclusions
Power grids are essential in combating climate change, requiring a transition to low-carbon energy and enhanced resilience against climate-induced extremes. The integration of VRE sources introduces complexities and uncertainties in grid operations, posing significant challenges for human operators and traditional solvers. Our work aims to foster progress towards these challenges by introducing RL2Grid, a benchmark designed to bridge the gap between current grid management practices and RL research. RL2Grid provides a standardized interface for power grid environments, featuring common rewards, state spaces, action spaces, and safety constraints across a pre-designed set of diverse and complex grid tasks in order to provide a common ground for monitoring and promoting progress. We perform a comprehensive evaluation of the performance of popular baselines on RL2Grid tasks, including versions augmented with domain-informed heuristics aimed at improving performance and sample efficiency, and find that there is still significant room for improvement in the performance of these methods. RL2Grid aims to accelerate algorithmic innovation towards improving power grid operations amidst the evolving challenges posed by climate change.
@@ Appendix A Relationship of RL2Grid to L2RPN tasks and solutions
We remind that our work builds on Grid2Op to provide a benchmark with standardized tasks, state and action spaces, rewards, and a safe (constrained) formalization, as well as comprehensive evaluation of common baselines inspired by L2RPN. These are critical to provide a common basis for assessing advances in RL methods (Papoudakis et al.,, 2021) as well as to improve accessibility to RL practitioners who may have limited prior knowledge of power systems.
@@ Appendix B RL baselines
PPO (Schulman et al.,, 2017) and Lagrangian PPO (LagrPPO) (Stooke et al.,, 2020). A PPO agent uses its neural network to directly approximate a policy. The agent learns the policy parameters by simplifying the TRPO (Schulman et al.,, 2015) algorithm, using a computationally tractable clipped objective. This clipping mechanism prevents large changes to the policy that could destabilize the training. At a high level, such a surrogate objective balances policy improvement and limits the divergence between policy updates. To drive the policy training, PPO also learns an advantage function to determine how much better (or worse) taking an action is compared to the expected value. By employing different probability distributions as a policy, a PPO agent can deal with both continuous (redispatching) and discrete (topological) actions. The Lagrangian version applies the same intuitions while learning additional value functions for each constraint. It then changes to policy training by considering the Lagrangian discussed in Section 2. In more detail, Lagrangian algorithms take gradient ascent steps in $\pi$ and descent steps in $\bm{\lambda}$ to trade off safety and task performance. These methods focus on satisfying the constraints using penalties $\bm{\lambda}$ that grows unbounded when constraints are violated. When constraints are satisfied, $\bm{\lambda}$ scale down (to zero), allowing the algorithm to maximize the task objective.
@@ Appendix D State Space
For the interested RL practitioner, we refer to the original Grid2Op documentation for exhaustive descriptions of these features (RTE France,, 2020).
@@ Appendix F Hyperparameters
Shared N° parallel envs 10, 20, 50 50
@@
$\alpha$ 0.1, 0.5, 1.0 1.0
$\beta$ 0.1, 0.5, 1.0 0.5
$\eta$ 0.1, 0.25, 0.5 0.5
$\lambda$ 0, 50 0 (TLO), 50 (LSI)
DQN Train frequency 50, 100, 1000 50
Target network update 1000, 5000, 10000 5000
Buffer size 100000, 250000, 500000, 1000000 1000000
@@
Entropy regularization 0.02, 0.2, 0.4 0.2
@@
Buffer size 100000, 250000, 500000, 1000000 1000000
Batch size 64, 128, 256 128
@@ Appendix G Omitted Figures in Section 5
The following figures report plots collected on Wandb (Wandb,, 2025). Figure G.1 shows the training curves for the remaining (discrete) topological action spaces. Due to the strict time limit imposed on the computation nodes (see Section 4) and the different computational requirements of the algorithms, not all the baselines perform the same number of steps in the time limit and the experiments with $36$ and $118$ bus consider 5 runs. The demands and limited performance of the topological baselines led us to exclude the results with the complete action space (i.e., difficulty set to 4). Additionally, despite the grid search of Table F.1, some baselines achieved lower performance than expected (e.g., SAC and DQN in the bus14 scenarios).
@@
Appendix H Omitted Performance Analysis in Section 5
Figure H.1: Average score for key operational components across heuristic and constraints-based baselines (reported on separate rows). The first column represents line margins, where higher values indicate better contingency management. The second column tracks overload penalties, with lower values reflecting improved grid stability. The third column captures topology modifications, showing the extent to which agents reconfigure the grid.